OpenStudy (anonymous):

What is common to the variation in the range and the height of a projectile? a - horizontal velocity b- time of flight vertical velocity, horizontal acceleration c - vertical velocity d - horizontal acceleration

sam (.sam.):

\[\Large v_y=u_y+at \\ \\ \Large v_y^2=u_y^2+2ah \\ \\ \Large h=u_yt+\frac{1}{2}at^2 \\ \\ \Large h=\frac{u_y+v_y}{2}t\] You can tell from these